# Towards Activated Muscle Group Estimation in the Wild

### v4

@@ Abstract.
In this paper, we tackle the new task of video-based Activated Muscle Group Estimation (AMGE) aiming at identifying active muscle regions during physical activity in the wild. To this intent, we provide the MuscleMap dataset featuring ${>}15K$ video clips with $135$ different activities and $20$ labeled muscle groups. This dataset opens the vistas to multiple video-based applications in sports and rehabilitation medicine under flexible environment constraints. The proposed MuscleMap dataset is constructed with YouTube videos, specifically targeting High-Intensity Interval Training (HIIT) physical exercise in the wild. To make the AMGE model applicable in real-life situations, it is crucial to ensure that the model can generalize well to numerous types of physical activities not present during training and involving new combinations of activated muscles. To achieve this, our benchmark also covers an evaluation setting where the model is exposed to activity types excluded from the training set. Our experiments reveal that the generalizability of existing architectures adapted for the AMGE task remains a challenge. Therefore, we also propose a new approach, TransM3E, which employs a multi-modality feature fusion mechanism between both the video transformer model and the skeleton-based graph convolution model with novel cross-modal knowledge distillation executed on multi-classification tokens. The proposed method surpasses all popular video classification models when dealing with both, previously seen and new types of physical activities. The contributed dataset and code will be publicly available at https://github.com/KPeng9510/MuscleMap.
Activate muscle group estimation, activity understanding, scene understanding.
@@ 1. Introduction
Human activity understanding is important as it enables the development of applications and systems that can enhance healthcare, improve security, and optimize various aspects of daily life by automatically identifying and understanding human actions and behaviors (Ahmad et al., 2021; Singh et al., 2023; Zhuang et al., 2023; Li et al., 2023; Jain et al., 2023; Subbareddy K and Devi L, 2023). Knowing which skeletal muscles of the human body are activated benefits human activity understanding, and sport and rehabilitation medicine from multiple perspectives and prevents inappropriate muscle usage which may cause physical injuries (Estember and Huang, 2019). In health care, patients need to know how to conduct the exercise correctly to recover from surgery (Kuiken et al., 2006) or specific diseases (Alibeji et al., 2018), e.g., COVID-19 (Rahma et al., 2020). Knowledge about muscle activations allows for user-centric fitness applications providing insights for everyday users or professional athletes who need specially adapted training. The majority of existing work on Activated Muscle Group Estimation (AMGE) is based on wearable devices with electrode sensors (de Vries et al., 2021). Yet, many wearable devices are inconvenient and heavy (Liang and Yuan, 2016), even harmful to health (Bababekova et al., 2011), and have limited usage time due to the battery (Seneviratne et al., 2017). A big strength of wearable devices is the high accuracy achieved through direct signal measurement from skin or muscle tissue. However, such exact bio-electrical changes are not required in a large number of medical recovery programs, and knowing the binary activation status of the muscle as shown in Figure 1 is sufficient in many situations (Williams et al., 2019; Sanchez-Lorente et al., 2018; Lu et al., 2019). In contrast to wearable devices, most people have a video camera available at hand on their phone or laptop. Applying video-based AMGE on in-the-wild data collected by using smartphones or other widely available smart devices would allow for the application of such programs even without access to specialized hardware. Thereby, end-to-end video-based AMGE approaches are expected to be developed to prevent overburdens caused by wearable devices from both physical and psychological points of view. Can modern deep learning algorithms relate fine-grained physical movements to individual muscles? To answer this question, we tackle the barely researched task of video-based active muscle group estimation under an in-the-wild setting, which estimates muscle contraction during physical activities from video recordings without a restricted environment and background constraints.
@@
We select various off-the-shelf Convolutional Neural Networks (CNNs) (Carreira and Zisserman, 2017; Feichtenhofer et al., 2019), Graph Convolutional Networks (GCNs) (Yan et al., 2018; Chen et al., 2021b; Lee et al., 2022b), and transformer-based architectures (Fan et al., 2021; Li et al., 2022; Liu et al., 2022c) from human activity recognition field together with statistic methods as baselines. Our proposed MuscleMap benchmark describes a multi-label classification problem where each sample might be annotated with one or up to twenty labels. However, we find that it is challenging for all these models when they deal with new activity types containing new activated muscle combinations at test time considering the AMGE generalizability. Skeleton-based models are observed to show good performance on the new activity types while working not well on known activity types. The video-based models show good performance on the known activity types while delivering limited performance on the new activity types. An approach that can work well on both known and new activity types is thereby expected.
@@
Knowledge distillation (Hinton et al., 2015) is leveraged for cross-modality knowledge transfer after the feature map reduction of the transformer block to enable a more informative latent space learning for different modalities. Transferring cross-modality knowledge during training significantly benefits the model in finding out cross-modality informative cues for the AMGE task. However, we find that it is difficult to achieve the knowledge distillation between two models with obvious architecture differences, e.g., graph convolutional networks (GCNs) and video transformers, considering the alignment of the feature maps coming from different backbones and modalities to achieve the appropriate and effective knowledge distillation. Aside from the architectural differences, we examine that using late fusion to fuse the skeleton-based model and video-based model can not achieve a satisfactory performance due to the lack of alignment of the two different feature domains.
We thereby propose a cross-modality MCT-level knowledge distillation scheme considering knowledge distillation on the intermediate and final layers by specifically designing a knowledge distillation MCT for the model of each modality. Alongside the MCT used solely for the classification, we leverage another MCT to execute the knowledge distillation for each modality. The cross-modality knowledge distillation is then executed only between the knowledge distillation MCTs from the two modalities, whereas existing works mostly use knowledge distillation calculated from the full embeddings or single knowledge distillation token at the final layer and use a larger teacher (Hinton et al., 2015; Monti et al., 2022; Fang et al., 2021; Touvron et al., 2021; Liu et al., 2022a). While the mentioned MCTKD mechanism integrates cross-modal knowledge into our main network with additional MCT for the knowledge distillation, another contribution, MCTF, merges the MCT of the distilled knowledge and the MCT for classification to achieve a final prediction towards the active muscle regions during human body motion for each modality. By combining these three components, TransM3E achieves state-of-the-art performances with superior generalizability compared to the tested baselines.
In summary, our contributions are listed as follows:
  • We open the vistas of video-based Activated Muscle Group Estimation in the wild task with the aim of lowering the threshold of entry to muscle-activation-based health care and sports applications.
  • We provide a new benchmark MuscleMap to propel research on the aforementioned task which includes the large-scale MuscleMap dataset. We also present a large number of baseline experiments for this benchmark, including CNN-, transformer-, and GCN-based approaches.
@@ 2. Related Work
Activity Recognition is a dominating field within visual human motion analysis (Ahmad et al., 2021; Singh et al., 2023; Zhuang et al., 2023; Li et al., 2023; Jain et al., 2023; Chen et al., 2021c; Yuan et al., 2020; Hao et al., 2020; Sun et al., 2022) which was propelled by the advent of Convolutional Neural Networks (CNNs) with 2D-CNNs (Feichtenhofer et al., 2016) in combination with recurrent neural networks (RNNS) (Donahue et al., 2015) or different variations of 3D-CNNs (Carreira and Zisserman, 2017; Feichtenhofer et al., 2019; Peng et al., 2022b). More recently, transformer-based methods advanced over 3D-CNNs, especially with advanced pre-training methods and large datasets (Li et al., 2022; Liu et al., 2022c, b). Action Quality Assessment (AQA) (Parmar and Morris, 2019; Tang et al., 2020) and Visual Calory Estimation (VCE) (Peng et al., 2022a) relate to our work since these methods likewise shift the question of research from what? to how? with the aim of detailed analysis of human motion. Multimodal data is a common strategy, e.g., by combining RGB video with audio (Piergiovanni et al., 2020; Patrick et al., 2020; Alayrac et al., 2020), poses (Schneider et al., 2022), optical flow (Piergiovanni et al., 2020), or temporal difference images (Panda et al., 2021). Skeleton data is also commonly used as a modality for activity recognition on their own. Yan et al.  (Yan et al., 2018) and follow-up research (Chen et al., 2021b; Lee et al., 2022b; Shi et al., 2019, 2020) make use of GCNs, while competitive approaches leverage CNNs with special pre-processing methods (Choutas et al., 2018; Duan et al., 2022).
@@
Datasets which combine visual data of the human body with muscle activation information are sparse and mainly limited to specific sub-regions of the human body, e.g., for hand gesture recognition (Gao et al., 2021). In contrast, a large variety of full-body human activity recognition datasets were collected in recent years, which are labelled with high-level human activities (Kuehne et al., 2011; Peng et al., 2023), fine grained human action segments (Zhao et al., 2019; Li et al., 2020), or action quality annotations (Shao et al., 2020). We leverage such datasets by extending them with muscle group activation labels.
@@ 3. Benchmark
Figure 2. An overview of the number of samples and the number of activity types per muscle region (@R), depicted at the top left and the top right. On the bottom, some activity-specific samples from the MuscleMap dataset are shown according to the corresponding muscle group.
@@ 3.1. MuscleMap Dataset
With the new video-based active muscle group estimation in-the-wild task in mind, we collect the MuscleMap dataset by querying YouTube for the physical exercise video series. The collected dataset contains $135$ activity types as well as $15,004$ video clips and is competitive compared to other video-based datasets targeting fine-grained tasks, as shown in Table 1. Twenty activities are reserved for the validation and test splits of new activities, which are not included in the training set. MuscleMap targets physical exercise videos from fitness enthusiasts. High-Intensity Interval Training (HIIT) exercises are well suited for the AMGE in-the-wild task since they display a large range of motions that are designed to activate specific muscle groups and instructional videos provide high-quality examples of the displayed motion. The collected videos in our dataset are mostly near-person, which can benefit video-based muscle contribution understanding for the in-the-wild videos. A small set of activities from the MuscleMap dataset is shown in the bottom part of Figure 2. In Table 1, MuscleMap is compared with existing human activity recognition, action quality assessment, calorie consumption datasets, and time series-wise muscle activation regression dataset.
@@ 3.3. Evaluation Protocol
Figure 3. TransM3E. The knowledge is distilled between the skeleton branch and the RGB branch, PTs denote the patch embedding tokens. Three main components are shown, e.g., Multi-Classification Tokens (MCTs), Multi-Classification Tokens Knowledge Distillation (MCTKD), and Multi-Classification Tokens Fusion (MCTF).
@@ 4.1. Preliminaries of MViT
TransM3E is based on the improved multi-scale visual transformer (MViTv2) (Li et al., 2022), which is based on MViTv1 (Fan et al., 2021). The model architecture of TransM3E is shown in Figure 3. Compared with ViT (Dosovitskiy et al., 2021), MViTv1 increases the channel resolution progressively and reduces the resolution on the spatiotemporal plane simultaneously, which realizes pooling operations both on Keys ($\mathbf{K}$) and Queries ($\mathbf{Q}$). The basic idea of MViTv1 is the construction of different low- and high-level visual modeling stages (Fan et al., 2021). Multi-scale pooling attention is one of the major components of MViTv2 compared with ViT. MViTv2 uses decomposed relative position embeddings and residual pooling connections to integrate the principle of shift-invariance into the model and reduce computational complexity, while the downscaling in MViTv1 is achieved by large strides on the Keys ($\mathbf{K}$) and Values ($\mathbf{V}$).
@@ 4.3. Multi-Classification Tokens Knowledge Distillation (MCTKD)
Multi-Classification Tokens Knowledge Distillation (MCTKD) is one of our main contributions. To the best of our knowledge, we are the first to introduce this technique which can enable knowledge distillation on the multi-classification tokens between two backbones with obvious structure differences. Through our observation in this work, we find that directly merging the feature from the skeleton-based model and video-based model can not achieve a satisfactory performance due to the huge structure and modality difference. To achieve appropriate feature fusion between two architectures on different modalities with obvious differences, we need a new solution focusing on this issue. Knowledge distillation, which has the capability to accomplish the feature space alignment, is firstly explored in our work on the MCT perspective to assist the cross-modality feature fusion for the AMGE in-the-wild task.
@@ 4.4. Multi-Classification Tokens Fusion (MCTF)
(8) $\mathbf{A}_{m}=$ $Sum(\mathbf{A}_{mm}^{m},\mathbf{A}_{mr}^{m},\mathbf{A}_{rm}^{m}),$
$\mathbf{A}_{r}=$ $Sum(\mathbf{A}_{rr}^{r},\mathbf{A}_{mr}^{r},\mathbf{A}_{rm}^{r}).$
@@
Table 2. Experimental results on the MuscleMap benchmark. Here, known val, new val, known test, and new test denote evaluation and test sets for normal and generalizable validation and test, respectively. mean val and mean test denote the averaged mean average precision (mAP) of normal and generalizable settings.
@@ 4.6. Analysis on the MuscleMap Benchmark
The results of different architectures on our benchmark are provided in Table 2. First, the approaches include Random, in which the muscle activation is predicted randomly, and All Ones, in which all the samples are predicted as using all the muscle regions. These two simple approaches are used to serve as statistic baselines. Random and All Ones show overall low performances with ${<}30\%$ mAP on all the evaluations. These statistical approaches are leveraged to make comparisons between deep-learning-based approaches to verify whether the model predicts muscle activation randomly or not. The skeleton-based approach, e.g., HD-GCN (Lee et al., 2022b), ST-GCN (Yan et al., 2018), and CTR-GCN (Chen et al., 2021b), obviously outperform the statistic approaches and deliver promising performances when dealing with unseen activity types. Video-based approaches surpass statistic and skeleton baselines in terms of the AMGE of the known activities, where transformer-based approaches, e.g., MViTv2 S/B (Li et al., 2022) and VideoSwin S/B (Liu et al., 2022c), and CNN-based approaches, e.g., C2D (Feichtenhofer et al., 2017), I3D (Carreira and Zisserman, 2017), Slow (Feichtenhofer et al., 2019), SlowFast (Feichtenhofer et al., 2019), are leveraged. MViTv2-S shows good performance compared with the other methods due to the capability for reasoning long-term information and the multi-scale pooling setting which can extract informative cues from different abstract perspectives, especially with $79.6\%$, $79.7\%$ for mean val and mean test on the MuscleMap dataset. However, we find that skeleton-based approaches work well on the new activities while they can not deliver satisfactory results on the known activities. On the other hand, video-based approaches work well on the known activities while they can not provide promising AMGE results on the new activities. A good AMGE model is expected to work well in both of the scenarios.
@@
To simultaneously achieve good performances for both the new activities and the known activities, we would like to grasp the advantages both of the skeleton-based approaches and the video-based approaches. We proposed TransM3E, which achieves feature fusion and knowledge distillation by using multi-classification tokens (MCT) on both the video-based model and the skeleton-based model, by using the most outperforming backbones from the two different modalities, i.e., MViTv2-S and HD-GCN. This new proposed method incorporates knowledge distillation from the multi-classification token level and feature fusion from the multi-classification token level to harvest more underlying attributes of the body motion which can benefit the activated muscle group estimation task in the wild task. TransM3E surpasses all the others by large margins. TransM3E is a transformer-based approach due to the capability for long-term reasoning of visual transformers (Vaswani et al., 2017) since the AMGE should consider the activities at the global level, which requires long-term information reasoning. TransM3E has $64.1\%$, $97.8\%$, $64.2\%$, and $81.0\%$ mAP considering new val, known val, new test, and known test on our benchmark, while the generalizability to new activities is mostly highlighted. TransM3E outperforms MViTv2-S by $1.4\%$ and $1.3\%$ on the mean val and mean test, which especially works well for new val and new test as TransM3E surpasses MViTv2-S by $2.7\%$ and $2.8\%$. Our method achieves significant improvements in AMGE through three main strategies: enhanced attribute reasoning using MCT, effective knowledge exchange between video and skeleton models, and integration of distillation and classification MCTs for better information consolidation. Despite its strengths, the performance gap in AMGE between known and new activities highlights areas for improvement, particularly in reducing misclassifications and biases, with more insights to be shared in a forthcoming ablation study analysis.

### v5

# Towards Video-based Activated Muscle Group Estimation in the Wild

@@ Abstract.
In this paper, we tackle the new task of video-based Activated Muscle Group Estimation (AMGE) aiming at identifying active muscle regions during physical activity in the wild. To this intent, we provide the MuscleMap dataset featuring ${>}15K$ video clips with $135$ different activities and $20$ labeled muscle groups. This dataset opens the vistas to multiple video-based applications in sports and rehabilitation medicine under flexible environment constraints. The proposed MuscleMap dataset is constructed with YouTube videos, specifically targeting High-Intensity Interval Training (HIIT) physical exercise in the wild. To make the AMGE model applicable in real-life situations, it is crucial to ensure that the model can generalize well to numerous types of physical activities not present during training and involving new combinations of activated muscles. To achieve this, our benchmark also covers an evaluation setting where the model is exposed to activity types excluded from the training set. Our experiments reveal that the generalizability of existing architectures adapted for the AMGE task remains a challenge. Therefore, we also propose a new approach, TransM3E, which employs a multi-modality feature fusion mechanism between both the video transformer model and the skeleton-based graph convolution model with novel cross-modal knowledge distillation executed on multi-classification tokens. The proposed method surpasses all popular video classification models when dealing with both, previously seen and new types of physical activities. The database and code can be found at https://github.com/KPeng9510/MuscleMap.
Activate muscle group estimation, video-based human activity understanding, video-based scene understanding.
@@ 1. Introduction
Human activity understanding is important as it enables the development of applications and systems that can enhance healthcare, improve security, and optimize various aspects of daily life by automatically identifying and understanding human actions and behaviors (Ahmad et al., 2021; Jain et al., 2023; Subbareddy K and Devi L, 2023). Knowing which skeletal muscles of the human body are activated benefits human activity understanding, and sport and rehabilitation medicine from multiple perspectives and prevents inappropriate muscle usage which may cause physical injuries (Estember and Huang, 2019). In health care, patients need to know how to conduct the exercise correctly to recover from surgery (Kuiken et al., 2006) or specific diseases (Alibeji et al., 2018), e.g., COVID-19 (Rahma et al., 2020). Knowledge about muscle activations allows for user-centric fitness applications providing insights for everyday users or professional athletes who need specially adapted training. The majority of existing work on Activated Muscle Group Estimation (AMGE) is based on wearable devices with electrode sensors (de Vries et al., 2021). Yet, many wearable devices are inconvenient and heavy (Liang and Yuan, 2016), even harmful to health (Bababekova et al., 2011), and have limited usage time due to the battery (Seneviratne et al., 2017). A big strength of wearable devices is the high accuracy achieved through direct signal measurement from skin or muscle tissue. However, such exact bio-electrical changes are not required in a large number of medical recovery programs, and knowing the binary activation status of the muscle as shown in Figure 1 is sufficient in many situations (Williams et al., 2019; Sanchez-Lorente et al., 2018; Lu et al., 2019). In contrast to wearable devices, most people have a video camera available at hand on their phone or laptop. Applying video-based AMGE on in-the-wild data collected by using smartphones or other widely available smart devices would allow for the application of such programs even without access to specialized hardware. Thereby, end-to-end video-based AMGE approaches are expected to be developed to prevent overburdens caused by wearable devices from both physical and psychological points of view. Can modern deep learning algorithms relate fine-grained physical movements to individual muscles? To answer this question, we tackle the barely researched task of video-based active muscle group estimation under an in-the-wild setting, which estimates muscle contraction during physical activities from video recordings without a restricted environment and background constraints.
@@
We select various off-the-shelf Convolutional Neural Networks (CNNs) (Carreira and Zisserman, 2017; Feichtenhofer et al., 2019), Graph Convolutional Networks (GCNs) (Yan et al., 2018; Chen et al., 2021b; Lee et al., 2022b), and transformer-based architectures (Fan et al., 2021; Li et al., 2022; Liu et al., 2022c) from the human activity recognition field, along with statistical methods, as baselines. Our proposed MuscleMap benchmark addresses a multi-label classification problem where each sample may have one to twenty labels. These models struggle with new activity types featuring new muscle combinations at test time, impacting AMGE generalizability. Skeleton-based models perform well on new activity types but not on known ones, whereas video-based models excel on known types but perform poorly on new ones. An approach that performs well on both known and new activity types is needed.
@@
Knowledge distillation (Hinton et al., 2015) is leveraged for cross-modality knowledge transfer after the feature map reduction of the transformer block to enable a more informative latent space learning for different modalities. Transferring cross-modality knowledge during training significantly benefits the model in finding out cross-modality informative cues for the AMGE task. However, we find that it is difficult to achieve the knowledge distillation between two models with obvious architectural differences, e.g., GCNs and video transformers, considering the alignment of the feature maps coming from different backbones and modalities to achieve the appropriate and effective knowledge distillation. Aside from the architectural differences, we examine that using late fusion to fuse the skeleton-based model and video-based model can not achieve a satisfactory performance due to the lack of alignment of the two different feature domains.
We propose a cross-modality MCT-level knowledge distillation scheme considering intermediate and final layer distillation by designing a specific knowledge distillation MCT for each modality. Alongside the classification MCT, another MCT executes knowledge distillation for each modality. Cross-modality knowledge distillation occurs only between the knowledge distillation MCTs from the two modalities, unlike existing works that use full embeddings or a single token at the final layer with a larger teacher (Hinton et al., 2015; Liu et al., 2022a). Our MCTKD mechanism integrates cross-modal knowledge into the main network, while MCTF merges the distilled knowledge MCT and classification MCT for the final prediction of active muscle regions during human body motion. Combining these components, TransM3E achieves state-of-the-art performance with superior generalizability compared to tested baselines. In summary, our contributions are listed as follows:
  • We propose a new task of video-based Activated Muscle Group Estimation in the wild task with the aim of lowering the threshold of entry to muscle-activation-based health care and sports applications.
  • We provide a new benchmark MuscleMap to propel research on the aforementioned task which includes the large-scale MuscleMap dataset. We also present baseline experiments for this benchmark, including CNN-, transformer-, and GCN-based approaches.
@@ 2. Related Work
Activity Recognition is a dominating field within visual human motion analysis (Ahmad et al., 2021; Jain et al., 2023) which was propelled by the advent of Convolutional Neural Networks (CNNs) with 2D-CNNs (Feichtenhofer et al., 2016) in combination with recurrent neural networks (RNNS) (Donahue et al., 2015) or different variations of 3D-CNNs (Carreira and Zisserman, 2017; Feichtenhofer et al., 2019; Peng et al., 2022b). More recently, transformer-based methods advanced over 3D-CNNs, especially with advanced pre-training methods and large datasets (Li et al., 2022; Liu et al., 2022c, b; Peng et al., 2024a). Action Quality Assessment (AQA) (Parmar and Morris, 2019; Tang et al., 2020) and Visual Calory Estimation (VCE) (Peng et al., 2022a) relate to our work since these methods likewise shift the question of research from what? to how? with the aim of detailed analysis of human motion. Multimodal data is a common strategy, e.g., by combining RGB video with audio (Piergiovanni et al., 2020; Patrick et al., 2020; Alayrac et al., 2020), poses (Schneider et al., 2022), optical flow (Piergiovanni et al., 2020), or temporal difference images (Panda et al., 2021). Skeleton data is also commonly used as a modality for activity recognition on its own. Yan et al.  (Yan et al., 2018) and follow-up research (Shi et al., 2020; Peng et al., 2024b; Xu et al., 2024; Wei et al., 2024) make use of GCNs, while competitive approaches leverage CNNs with special pre-processing methods (Choutas et al., 2018; Duan et al., 2022).
@@
Datasets which combine visual data of the human body with muscle activation information is sparse and mainly limited to specific sub-regions of the human body, e.g., for hand gesture recognition (Gao et al., 2021). In contrast, a large variety of full-body human activity recognition datasets were collected in recent years, which are labeled with high-level human activities (Kuehne et al., 2011; Peng et al., 2023), fine-grained human action segments (Zhao et al., 2019; Li et al., 2020), or action quality annotations (Shao et al., 2020). We leverage such datasets by extending them with muscle group activation labels.
@@ 3. Benchmark
Figure 2. An overview of the number of samples and the number of activity types per muscle region (@R), depicted at the top left and the top right.
@@ 3.1. MuscleMap Dataset
With the new video-based active muscle group estimation in-the-wild task in mind, we collect the MuscleMap dataset by querying YouTube for the physical exercise video series. The collected dataset contains $135$ activity types as well as $15,004$ video clips and is competitive compared to other video-based datasets targeting fine-grained tasks, as shown in Table 1. Twenty activities are reserved for the validation and test splits of new activities, which are not included in the training set. MuscleMap targets physical exercise videos from fitness enthusiasts. High-Intensity Interval Training (HIIT) exercises are well suited for the AMGE in-the-wild task since they display a large range of motions that are designed to activate specific muscle groups and instructional videos provide high-quality examples of the displayed motion. The collected videos in our dataset are mostly near-person, which can benefit video-based muscle contribution understanding for the in-the-wild videos. To deliver diverse modalities, we provide RGB, RGB Diff, optical flow extracted by DenseFlow (Wang et al., 2020), and 2D skeleton data extracted through AlphaPose (Fang et al., 2023). A small set of activities from the MuscleMap dataset is shown in the bottom part of Figure 2. In Table 1, MuscleMap is compared with existing human activity recognition, action quality assessment, calorie consumption datasets, and time series-wise muscle activation regression dataset. We ensured that all YouTube videos used were publicly available and complied with the platform’s terms of service.
@@ 3.3. Evaluation Protocol
Figure 3. An overview of the proposed TransM3E architecture.
@@ 4.1. Preliminaries of MViT
TransM3E is based on MViTv2 (Li et al., 2022). The model architecture of TransM3E is shown in Figure 3. MViTv2 uses decomposed relative position embeddings and residual pooling connections to integrate shift-invariance and reduce computational complexity, while MViTv1 achieves downscaling by large strides on Keys ($\mathbf{K}$) and Values ($\mathbf{V}$).
@@ 4.3. Multi-Classification Tokens Knowledge Distillation (MCTKD)
Multi-Classification Tokens Knowledge Distillation (MCTKD) is one of our main contributions. We are the first to introduce this technique, enabling knowledge distillation on multi-classification tokens between two structurally different backbones. Directly merging features from skeleton-based and video-based models underperforms due to structural and modality differences. To achieve effective feature fusion between these distinct architectures, we need a new solution. Our work explores knowledge distillation for feature space alignment from the MCT perspective, aiding cross-modality feature fusion for the AMGE in-the-wild task.
@@ 4.4. Multi-Classification Tokens Fusion (MCTF)
(8) $\mathbf{A}_{m},\mathbf{A}_{r}=Sum(\mathbf{A}_{mm}^{m},\mathbf{A}_{mr}^{m},\mathbf{A}_{rm}^{m}),Sum(\mathbf{A}_{rr}^{r},\mathbf{A}_{mr}^{r},\mathbf{A}_{rm}^{r}).$
@@
Table 2. Experimental results on the MuscleMap benchmark.
@@ 5.2. Analysis on the MuscleMap Benchmark
The results of different architectures on our benchmark are provided in Table 2. First, the approaches include Random, in which the muscle activation is predicted randomly, and All Ones, in which all the samples are predicted as using all the muscle regions. These two simple approaches are used to serve as statistic baselines. Random and All Ones show overall low performances with ${<}30\%$ mAP on all the evaluations. These statistical approaches are leveraged to make comparisons between deep-learning-based approaches to verify whether the model predicts muscle activation randomly or not. The skeleton-based approach, e.g., HD-GCN (Lee et al., 2022b), ST-GCN (Yan et al., 2018), and CTR-GCN (Chen et al., 2021b), obviously outperform the statistic approaches and deliver promising performances when dealing with unseen activity types. Video-based approaches surpass statistic and skeleton baselines in terms of the AMGE of the known activities, where transformer-based approaches, e.g., MViTv2 S/B (Li et al., 2022) and VideoSwin S/B (Liu et al., 2022c), and CNN-based approaches, e.g., C2D (Feichtenhofer et al., 2017), I3D (Carreira and Zisserman, 2017), Slow (Feichtenhofer et al., 2019), SlowFast (Feichtenhofer et al., 2019), are leveraged. MViTv2-S shows good performance due to its ability to reason long-term information and its multi-scale pooling setting, achieving $79.6\%$ and $79.7\%$ for mean val and mean test on the MuscleMap dataset. However, skeleton-based approaches perform well on new activities but not on known ones due to the lack of visual appearance, while video-based approaches excel on known activities but not on new ones due to the sensitivity to the background changes. A good AMGE model should perform well in both scenarios.
@@
To achieve this, we propose TransM3E, which combines the advantages of both skeleton-based and video-based approaches. It uses multi-classification tokens (MCT) for feature fusion and knowledge distillation, leveraging the top-performing backbones from both modalities: MViTv2-S and HD-GCN. TransM3E surpasses all the others by large margins. TransM3E is a transformer-based approach due to the capability for long-term reasoning of visual transformers (Vaswani et al., 2017) since the AMGE should consider the activities at the global level, which requires long-term information reasoning. TransM3E has $64.1\%$, $97.8\%$, $64.2\%$, and $81.0\%$ mAP considering new val, known val, new test, and known test on our benchmark, while the generalizability to new activities is mostly highlighted. TransM3E outperforms MViTv2-S by $1.4\%$ and $1.3\%$ on the mean val and mean test, which especially works well for new val and new test as TransM3E surpasses MViTv2-S by $2.7\%$ and $2.8\%$. During the experiments, we observe that the obliques group is the hardest region to achieve AMGE. We also conduct per-label analysis towards sports with body weights and find that the AMGE performance of the motions with fitness equipment is higher than those with body weight.